OpenStudy (anonymous):

Cameron is replacing the roof on a gazebo. The roof is in the shape of a cone with an open bottom. The diameter is fourteen feet and the slant height is ten feet. What is the lateral area of the roof? Use three point one four for pi. one hundred sixty-three point three square feet two hundred nineteen point eight square feet four hundred thirty-nine point six square feet five hundred twelve point nine square feet

OpenStudy (igreen):

Hold on..

OpenStudy (igreen):

The formula is \(\pi r h\). Where: \(\pi\) = 3.14 r = Radius h = Height

OpenStudy (igreen):

\(\pi r h\) \((3.14)(7)(10)\) (It gives you the diameter, and the radius is half the diameter which gives us 7) \((21.98)(10)\) \(219.8\)

OpenStudy (igreen):

@Gravity_Dreams

OpenStudy (anonymous):

thanks